assume a vachile's engine is producing 300 ft.-lb. of torque at 4000 rpm. If the vechile's four-speed manual transmission is in third gear (with ratio 1.5:1),then:
(a) The transmission output shaft speed is about ____rpm.
(b) The transmission output shaft torque is about ___ ft.-lb.

Answer by Alan3354(69443)

assume a vachile's [sic] engine is producing 300 ft.-lb. of torque at 4000 rpm. If the vechile's [sic] four-speed manual transmission is in third gear (with ratio 1.5:1),then:
(a) The transmission output shaft speed is about ____rpm.
1.5/1 = 4000/w
1.5w = 4000
w = 8000/3 rpm
-----------
(b) The transmission output shaft torque is about ___ ft.-lb.
Torque = 300*1.5 = 450 ft-lb